OpenStudy (cookiemonster18):

suppose y varies directly with X write a direct variation equation that relates X to Y Y=10.4 when X=4 A.y=3.4x B.y=3.8x C.y=3.2x D.y=2.6x (I think it's D am I correct?)

OpenStudy (anonymous):

is tis the equation to the problem

OpenStudy (anonymous):

You are correct, the answer is D. If you plug in x=4 you find y = 2.6 * 4 = 10.4, which is what we expect.

OpenStudy (danjs):

that is the easy way to do it if they give you answer choices all in the form of a direct variation y=k*x

OpenStudy (danjs):

It is a line through the origin (0,0)